The intensity of radiation reaching the earth from the sun is 1350 W/m2. The earth's radius is 6.4 ´ 106 m. How big a force does this radiation exert on the earth? (Assume it is all absorbed.)
Answer #1

Radiation pressure for absorption is given by

absorb cos Q

Ef = 1350

c=2.998 m/s

Considering the projected area we can assume \alpha = 0

This pressure acts on a projected area of

A=\pi (Re)^{2}= 3.14159 * (6.4*10^6)^2 =1.2867* 10^4 m^2

P=\frac{Ef}{c}=4.503*10^{-6} N/m^2

Radiation force = P*A

F=P*A=4.503* 10^{-6}*1.286796351*10^{14} =579.44*10^6 Newtons

In other words

59066 tons (that's huge)

Reference:

http://en.wikipedia.org/wiki/Radiation_pressure